# Impact of 3D macro physics and nuclear physics on the $p$ nuclei in $\mathrm{O}$-$\mathrm{C}$ shell mergers

### v1

@@ 2.1 Initial Model and Post-Processing Setup
where $c$ is the maximum extent of the dip, $w$ is the width, and $a$ is the center of the dip. The GOSH-like convective splitting could occur at the location where $D_{\alpha}=1$ for a significant burning event (Herwig et al., 2011). We centre this event at $a=4.95\;\mathrm{\mathrm{M}\mathrm{m}}$ at a location of probable energetic feedback could occur. The $\mathrm{O}$-shell could partially merge with the $\mathrm{C}$-shell due to feedback effects, so we consider a partial merger at $a=7.5\;\mathrm{\mathrm{M}\mathrm{m}}$ where the unmerged MLT profile has a dip as seen in Figure 3. A width of $w=0.25\;\mathrm{\mathrm{M}\mathrm{m}}$ is used for both scenarios, which is approximately the distance between the top of the O-shell and the convective bump above it in Figure 3. For both scenarios we consider a weaker dip to $c=10^{14}\;\mathrm{\mathrm{c}\mathrm{m}^{2}\mathrm{s}^{-1}}$ and a stronger dip to $c=10^{13}\;\mathrm{\mathrm{c}\mathrm{m}^{2}\mathrm{s}^{-1}}$ to investigate the impact on nucleosynthesis. The profiles for the GOSH-like and partial merger scenarios are shown in Figure 5.
@@
2.3 Determining impact of varying nuclear reactions
Many unstable $n$-deficient isotopes from $\mathrm{Se}$ to $\mathrm{Po}$ have not been measured experimentally and are determined by theoretical models, and the uncertainties of unmeasured reactions for unstable isotopes are much greater than for stable isotopes. To determine the impact of nuclear physics for the $\gamma$ process, we vary our $(\gamma,\alpha)$, $(\gamma,p)$, and $(\gamma,n)$ photo-disintegration rates used by the NuGrid code (Pignatari et al., 2016) for all unstable $n$-deficient isotopes from $\mathrm{Se}$ to $\mathrm{Po}$ by a random factor uniformly selected between 0.1 to 10 by a Monte Carlo method for 1000 cases. This applies the same approach used for $(n,\gamma)$ rates during the $i$ process developed by Denissenkov et al. (2018) and Denissenkov et al. (2021). This was done for the MLT and downturn mixing scenarios in Figure 4 with an ingestion rate of $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$. This approach also allows for the identification of reaction rates that are relevant for the production of an isotope using correlations. The Pearson coefficient describes correlations between $X/X_{\mathrm{no~{}variation}}$ and the variation factors where $X$ is the final mass-averaged and decayed mass fraction for a Monte Carlo case and $X_{\mathrm{no~{}variation}}$ is the same for the default case where all variation factors are $1$. All correlations with $|r_{\mathrm{P}}|\geq 0.15$ are reported in this study. In addition to the Pearson coefficient, a logarithmic slope $\zeta$ is also reported to determine the importance of a reaction on the final mass fraction of an isotope, which is discussed along with caveats about correlation rates in Appendix A.
3 Results
@@
MLT $-0.11$ $1.06$ $1.92$ $2.24$ 0.56
1×1$\times$1 × Downturn $0.05$ $1.12$ $1.98$ $2.58$ 0.59
3×3$\times$3 × Downturn $-0.23$ $1.28$ $2.18$ $2.83$ 0.69
10×10$\times$10 × Downturn $-1.23$ $1.18$ $2.10$ $2.89$ 0.76
50×50$\times$50 × Downturn $-5.47$ $0.88$ $1.81$ $2.72$ 0.79
@@
3.1 Convective-reactive production of the $p$ nuclei
[…]
Figure 6: Mass fractions without radiogenic contributions at $t=110~{}\mathrm{sec}$ for the MLT mixing scenario without $\mathrm{C}$-shell ingestion.
Locations of peak destruction can be identified by a sudden drop in mass fraction as seen for $\mathrm{{}^{156}Dy}$, $\mathrm{{}^{196}Hg}$, and to a lesser extent $\mathrm{{}^{144}Sm}$. This is the location where $D_{\alpha}=1$, which is different for each isotope (Herwig et al., 2011). This is contrary to the normal assumption of a well-mixed convective environment where $D_{\alpha}\ll 1$ or radiative burning where little to no mixing occurs, and is a key reason why the $\gamma$ process in the $\mathrm{O}$-shell produces the whole range of $p$ nuclei. Figure 7 shows the dominant reactions that $\mathrm{{}^{156}Dy}$ is produced and destroyed by in the simulation without $\mathrm{C}$-shell ingestion. These reactions are shown as reaction fluxes $f_{ij}$ which are defined as the net flux of a reaction between species $i$ and $j$:
@@ 3.1 Convective-reactive production of the $p$ nuclei
Figure 7: Reaction fluxes $f_{ij}$ for $\mathrm{{}^{156}Dy}$ and mass fractions $X_{i}$ for $\mathrm{{}^{156-158}Dy}$ without radiogenic contributions for the MLT scenario with no ingestion after $t=110\;\mathrm{\mathrm{s}}$. The direction of reactions $i\rightarrow j$ are written as left to right in the legend.
The reactions that $\mathrm{{}^{156}Dy}$ undergoes in the $\mathrm{O}$-shell depend on the location in the shell, however Figure 7 shows that the dominant destruction channel $\mathrm{{}^{156}Dy}(\gamma,n)\mathrm{{}^{155}Dy}$ net destroys $\mathrm{{}^{156}Dy}$, as Figure 11 will show. It is also evident that the mass fraction is not well-mixed as the gradient of the mass fraction is steep at the location of peak destruction.
Entraining $\mathrm{C}$-shell material is important for this process, as heavier species can be gradually depleted by $(\gamma,\alpha)$ and $(\gamma,p)$ reactions. Figure 8 shows the same as Figure 7, but with a $\mathrm{C}$-shell ingestion rate of $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, the maximum considered in this study. Since the initial amount of $\mathrm{{}^{156}Dy}$ in the ingested $\mathrm{C}$-shell is negligible as shown in Figure 2, the merger provides other species that are used to produce $\mathrm{{}^{156}Dy}$.
Figure 8: Reaction fluxes $f_{ij}$ for $\mathrm{{}^{156}Dy}$ and mass fractions $X_{i}$ for $\mathrm{{}^{156-158}Dy}$ without radiogenic contributions for the MLT scenario with an ingestion rate of $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$ after $t=110\;\mathrm{\mathrm{s}}$. The direction of reactions $i\rightarrow j$ are written as left to right in the legend.
Figure 9: Chart of reactions between isotopes at $m=1.64\;\mathrm{\mathrm{M}_{\odot}}$ [$T=2.28\;\mathrm{\mathrm{G}\mathrm{K}}$] for the same conditions as Figure 8. Both arrow colour and size indicate $\log_{10}(f_{ij})$, and arrows point in the direction of the reaction. The range of $\log_{10}(f_{ij})$ is the same as Figure 8.
There are several differences between Figures 7 and 8. First, $f_{ij}$ and $X_{i}$ are larger by several orders of magnitude and $\mathrm{{}^{156}Dy}$ has a net production in the shell, as Figure 11 will show in Section 3.3. Second, the mass fraction of $\mathrm{{}^{156}Dy}$ has a tilt up where it is net produced and then drops off sharply at the location of peak destruction instead of the decline seen in Figure 7.
@@
[…]
@@
Figures 7 and 8 demonstrate that in a convective-reactive environment isotopes are not well-mixed, and that the relevance of a reaction rate depends on the location in the shell. Figure 8 also show how isotopes can contribute to the production of each other at different locations in the shell. Finally, comparing Figure 8 to Figure 7 shows the importance of ingesting $\mathrm{C}$-shell material for significant production of the $p$ nuclei in the $\mathrm{O}$-shell.
3.2 Impact from a downturn and boosting mixing speeds
@@
Here we present the impact of a 3D-inspired gradual downturn at the lower boundary and boosting mixing speeds as explained in Section 2.2 with the diffusion coefficients in Figure 4. These cases are calculated with an ingestion speed of $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$ and the results are shown in Figure 10.
@@
Figure 10: The overproduction compared to initial of the $p$ nuclei for the MLT and 3D-inspired mixing scenarios. The average spread in production $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=0.96\;\mathrm{\,\mathrm{dex}}$. $\mathrm{OP}=0$ is the initial amount.
@@
The MLT simulation has an $\langle\mathrm{OP}\rangle$ of $2.24\;\mathrm{\,\mathrm{dex}}$, and the downturn scenarios have $\langle\mathrm{OP}\rangle$ of $2.58\;\mathrm{\,\mathrm{dex}}$, $2.83\;\mathrm{\,\mathrm{dex}}$, $2.89\;\mathrm{\,\mathrm{dex}}$, and $2.72\;\mathrm{\,\mathrm{dex}}$ for the 1×1$\times$1 ×, 3×3$\times$3 ×, 10×10$\times$10 ×, and 50×50$\times$50 ×3D-inspired mixing scenarios respectively. The average spread in production for each isotope $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=0.96\;\mathrm{\,\mathrm{dex}}$, which shows that mixing speeds are important for the production of the $p$ nuclei. This $\mathrm{O}$-shell during the merger significantly produces all $p$ nuclei except $\mathrm{{}^{74}Se}$, $\mathrm{{}^{78}Kr}$, and $\mathrm{{}^{84}Sr}$.
@@
The 3D-inspired 1×1$\times$1 × scenario favours the production of the heavier $p$ nuclei compared to the MLT scenario because $\tau_{\mathrm{mix}}$ decreases as the temperature. Because of this, more reactions occur at the cooler temperatures where the $(n,\gamma)$ and $(\gamma,n)$ reactions are more favoured than the $(\gamma,\alpha)$ and $(\gamma,p)$ reactions. All isotopes are comparably produced to MLT or more produced except $\mathrm{{}^{74}Se}$, $\mathrm{{}^{78}Kr}$, and $\mathrm{{}^{84}Sr}$ who require the hottest temperatures for their production.
@@
As mixing speeds increase, the production increases in a non-linear and non-monotonic way. For the 50×50$\times$50 × case, the average production of all $p$ nuclei is lower than the 3×3$\times$3 × and 10×10$\times$10 × scenarios. This is because the mixing speeds are high enough that material is advected to the bottom of the $\mathrm{O}$-shell fast enough despite the downturn, and correspondingly the lighter $p$ nuclei are generally more favoured including $\mathrm{{}^{74}Se}$, $\mathrm{{}^{78}Kr}$, and $\mathrm{{}^{84}Sr}$. Production for individual isotopes also can be non-linear and non-monotonic. For example, $\mathrm{{}^{115}Sn}$, $\mathrm{{}^{132}Ba}$, and $\mathrm{{}^{138}La}$ all increase for the 1×1$\times$1 × and 3×3$\times$3 × scenarios, but then their production is not as strong for the 10×10$\times$10 × and 50×50$\times$50 × scenarios.
@@
Another result is that isotopic pairs of the same element are not affected the same way by the downturn compared to MLT, nor by the increase in mixing speed and we find that the ratio between these isotopes is dependent on the mixing scenario. Since these isotopes are connected by $(\gamma,n)$ and $(n,\gamma)$ reactions, if the location of $D_{\alpha}=1$ for a reaction changes because of a change to the mixing speed or the presence of the decreasing diffusion profile, then the production of these isotopes will change. $\mathrm{{}^{94}Mo}$ is produced more for all mixing scenarios except for the 50×50$\times$50 × scenario, where $\mathrm{{}^{92}Mo}$ has a larger $\mathrm{OP}$, and likewise $\mathrm{{}^{115}Sn}$ exhibits this behaviour when compared to $\mathrm{{}^{112}Sn}$ and $\mathrm{{}^{114}Sn}$. We find that it is possible that the ratio can tend to unity as mixing speeds increase as seen for $\mathrm{{}^{96,98}Ru}$, $\mathrm{{}^{106,108}Cd}$, and $\mathrm{{}^{112,114}Sn}$. For $\mathrm{{}^{136,138}Ce}$ the lighter isotope is always favoured as mixing speed increases and for $\mathrm{{}^{156,158}Dy}$ the opposite is true. This demonstrates the importance of the decreasing diffusion profile and increasing mixing speed is for the production of the $p$ nuclei.
3.3 Impact from varying the ingestion rate

### v2

@@ II.1 Initial Model and Post-Processing Setup
where $c$ is the maximum extent of the dip, $w$ is the width, and $a$ is the center of the dip. The GOSH-like convective splitting could occur at the location where $D_{\alpha}=1$ for a significant burning event (Herwig et al., 2011). We centre this event at $a=4.95\;\mathrm{\mathrm{M}\mathrm{m}}$, where $1\;\mathrm{\mathrm{M}\mathrm{m}}=10^{6}\;\mathrm{\mathrm{m}}$, at a location of probable energetic feedback could occur. The $\mathrm{O}$ shell could partially merge with the $\mathrm{C}$ shell due to feedback effects, so we consider a partial merger at $a=7.5\;\mathrm{\mathrm{M}\mathrm{m}}$ where the unmerged MLT profile has a dip as seen in Figure 3. A width of $w=0.25\;\mathrm{\mathrm{M}\mathrm{m}}$ is used for both scenarios, which is approximately the distance between the top of the $\mathrm{O}$ shell and the convective bump above it in Figure 3. For both scenarios we consider a weaker dip to $c=10^{14}\;\mathrm{\mathrm{c}\mathrm{m}^{2}\mathrm{s}^{-1}}$ and a stronger dip to $c=10^{13}\;\mathrm{\mathrm{c}\mathrm{m}^{2}\mathrm{s}^{-1}}$ to investigate the impact on nucleosynthesis. The profiles for the GOSH-like and partial merger scenarios are shown in Figure 5.
@@
II.3 Determining impact of varying nuclear reactions
Many of the reactions involving the unstable $n$-deficient isotopes from $\mathrm{Se}$ to $\mathrm{Po}$ have not been measured experimentally and are determined by theoretical models, and the uncertainties of unmeasured reactions for unstable isotopes are much greater than for stable isotopes. To determine the impact of nuclear physics for the $\gamma$ process, we vary our $(\gamma,\alpha)$, $(\gamma,p)$, and $(\gamma,n)$ photo-disintegration rates used by the NuGrid code (Pignatari et al., 2016) for all unstable $n$-deficient isotopes from $\mathrm{Se}$ to $\mathrm{Po}$ by a random factor uniformly selected between 0.1 to 10 by a Monte Carlo method for 1000 cases. This applies the same approach used for $(n,\gamma)$ rates during the $i$ process developed by Denissenkov et al. (2018) and Denissenkov et al. (2021). This was done for the MLT and downturn mixing scenarios in Figure 4 with an ingestion rate of $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$. This approach also allows for the identification of reaction rates that are relevant for the production of an isotope using correlations. The Pearson coefficient describes correlations between $X/X_{\mathrm{no~variation}}$ and the variation factors where $X$ is the final mass-averaged and decayed mass fraction for a Monte Carlo case and $X_{\mathrm{no~variation}}$ is the same for the default case where all variation factors are $1$. All correlations with $|r_{\mathrm{P}}|\geq 0.15$ are reported in this study. In addition to the Pearson coefficient, a logarithmic slope $\zeta$ is also reported to determine the importance of a reaction on the final mass fraction of an isotope, which is discussed along with caveats about correlation rates in Appendix A.
III Results
@@
MLT 0.11-0.11 $1.06$ $1.92$ $2.24$ $0.56$
$1\times$ Downturn $0.05$ $1.12$ $1.98$ $2.58$ $0.59$
$3\times$ Downturn 0.23-0.23 $1.28$ $2.18$ $2.83$ $0.69$
$10\times$ Downturn 1.23-1.23 $1.18$ $2.10$ $2.89$ $0.76$
$50\times$ Downturn 5.47-5.47 $0.88$ $1.81$ $2.72$ $0.79$
@@
III.1 Convective-reactive production of the $p$ nuclei
Convective-reactive nucleosynthesis is characterized by a region where the timescales for advection and nuclear reactions are similar. In the $\gamma$ process, heavier species are produced at cooler temperatures of $2.3{-}2.5\;\mathrm{\mathrm{G}\mathrm{K}}$ and destroyed at higher temperatures, and lighter species are produced in temperatures up to $3.5\;\mathrm{\mathrm{G}\mathrm{K}}$ (Rauscher et al., 2013). Whether an isotope undergoes $(\gamma,n)$ and $(n,\gamma)$ reactions or contributes to the production of lighter $p$ nuclei by $(\gamma,p)$ and $(\gamma,\alpha)$ reactions depends on the temperature at that position. The convective-reactive environment of the $\mathrm{O}$ shell allows for the production of both light and heavy $p$ nuclei because the shell is not well-mixed, which allows the shell to produce most of the $p$ nuclei, although temperatures in the $\mathrm{O}$ shell are too cool to sufficiently produce those with $A<110$. Figure 6 shows how different mass ranges of $p$ nuclei can be produced and peak at different positions in the $\mathrm{O}$ shell.
Figure 6: Undecayed mass fractions at $t=110~\mathrm{sec}$ for the MLT mixing scenario without $\mathrm{C}$-shell ingestion.
Locations of peak destruction can be identified by a sudden drop in mass fraction as seen for $\mathrm{{}^{156}Dy}$, $\mathrm{{}^{196}Hg}$, and to a lesser extent $\mathrm{{}^{144}Sm}$. This is the location where $D_{\alpha}=1$, which is different for each isotope (Herwig et al., 2011). This is contrary to the normal assumption of a well-mixed convective environment where $D_{\alpha}\ll 1$ or radiative burning where little to no mixing occurs, and is a key reason why the $\gamma$ process in the $\mathrm{O}$ shell produces the whole range of $p$ nuclei. Figure 7 shows the dominant reactions that $\mathrm{{}^{156}Dy}$ is produced and destroyed by in the simulation without $\mathrm{C}$-shell ingestion. These reactions are shown as reaction fluxes $f_{ij}$ which are defined as the net flux of a reaction between species $i$ and $j$:
@@ III.1 Convective-reactive production of the $p$ nuclei
Figure 7: Reaction fluxes $f_{ij}$ for $\mathrm{{}^{156}Dy}$ and undecayed mass fractions $X_{i}$ for $\mathrm{{}^{156-158}Dy}$ for the MLT scenario with no ingestion after $t=110\;\mathrm{\mathrm{s}}$. Reactions in the legend are written in the direction $i\rightarrow j$
The reactions that $\mathrm{{}^{156}Dy}$ undergoes in the $\mathrm{O}$ shell depend on the location in the shell, however Figure 7 shows that the dominant destruction channel $\mathrm{{}^{156}Dy}(\gamma,n)\mathrm{{}^{155}Dy}$ net destroys $\mathrm{{}^{156}Dy}$, as Figure 11 will show. It is also evident that the mass fraction is not well-mixed as the gradient of the mass fraction is steep at the location of peak destruction.
Entraining $\mathrm{C}$-shell material is important for this process, as heavier species can be gradually depleted by $(\gamma,\alpha)$ and $(\gamma,p)$ reactions. Figure 8 shows the same as Figure 7, but with a $\mathrm{C}$-shell ingestion rate of $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, the maximum considered in this study. Since the initial amount of $\mathrm{{}^{156}Dy}$ in the ingested $\mathrm{C}$ shell is negligible as shown in Figure 2, the role of the merger is to provide species in the $\mathrm{C}$ shell that are critical for the production of $\mathrm{{}^{156}Dy}$ such as the stable $\mathrm{Dy}$ isotopes that undergo a $(\gamma,n)$ photodisintegration chain and the stable $\mathrm{Er}$ isotopes that do a sequence of $(\gamma,n)$ until the $\mathrm{{}^{160}Er}(\gamma,\alpha)\mathrm{{}^{156}Dy}$.
Figure 8: Reaction fluxes $f_{ij}$ for $\mathrm{{}^{156}Dy}$ and undecayed mass fractions $X_{i}$ for $\mathrm{{}^{156-158}Dy}$ for the MLT scenario with an ingestion rate of $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$ after $t=110\;\mathrm{\mathrm{s}}$. Reactions in the legend are written in the direction $i\rightarrow j$.
Figure 9: Chart of reactions between isotopes at $m=1.64\;\mathrm{\mathrm{M}_{\odot}}$ $T=2.28\;\mathrm{\mathrm{G}\mathrm{K}}$] for the same conditions as Figure 8. Both arrow colour and size indicate $\log_{10}(f_{ij})$, and arrows point in the direction of the reaction. The range of $\log_{10}(f_{ij})$ is the same as Figure 8.
There are several differences between Figures 7 and 8. First, $f_{ij}$ and $X_{i}$ are larger by several orders of magnitude and $\mathrm{{}^{156}Dy}$ has a net production in the shell, as Figure 11 will show in Section III.3. Second, the mass fraction of $\mathrm{{}^{156}Dy}$ has a tilt up where it is net produced and then drops off sharply at the location of peak destruction instead of the decline seen in Figure 7.
@@
Figure 8 also shows another feature of this convective-reactive environment: $\mathrm{{}^{156}Dy}$ and $\mathrm{{}^{157}Dy}$ co-produce each other. $\mathrm{{}^{156}Dy}$ is advected from its location of peak production at $\sim$$1.63\;\mathrm{\mathrm{M}_{\odot}}$ both deeper into the shell where it is fully destroyed, and toward the top where it undergoes $\mathrm{{}^{156}Dy}(n,\gamma)\mathrm{{}^{157}Dy}$, which mildly contributes to the production of $\mathrm{{}^{157}Dy}$ and peaks at $\sim$$1.73\;\mathrm{\mathrm{M}_{\odot}}$.
@@
Figures 7 and 8 demonstrate that in a convective-reactive environment isotopes are not well-mixed, and that the relevance of a reaction rate depends on their location in the shell. Figure 8 also show how isotopes can contribute to the production of each other at different locations in the shell. Finally, comparing Figure 8 to Figure 7 shows the importance of ingesting $\mathrm{C}$-shell material for significant production of the $p$ nuclei in the $\mathrm{O}$ shell.
III.2 Impact from a downturn and boosting mixing speeds
@@
Here we present the impact of a 3D-inspired gradual downturn at the lower boundary and boosting mixing speeds as explained in Section II.2 with the diffusion coefficients in Figure 4. These cases are calculated with an ingestion rate of $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$ and the results are shown in Figure 10.
@@
Figure 10: The overproduction compared to initial of the $p$ nuclei for the MLT and 3D-inspired mixing scenarios. The average spread in production $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=0.96\;\mathrm{\,\mathrm{dex}}$. $\mathrm{OP}=0\,\mathrm{dex}$ is the initial amount. Average $\mathrm{OP}$ for each scenario are provided as dashed lines and corresponds to the values presented in Table 1.
@@
The MLT simulation has an $\langle\mathrm{OP}\rangle$ of $2.24\;\mathrm{\,\mathrm{dex}}$, and the downturn scenarios have $\langle\mathrm{OP}\rangle$ of $2.58\;\mathrm{\,\mathrm{dex}}$, $2.83\;\mathrm{\,\mathrm{dex}}$, $2.89\;\mathrm{\,\mathrm{dex}}$, and $2.72\;\mathrm{\,\mathrm{dex}}$ for the $1\times$, $3\times$, $10\times$, and $50\times$3D-inspired mixing scenarios respectively. The average spread in production for each isotope $\mathrm{OP}_{\max}-\mathrm{OP}_{\min}=0.96\;\mathrm{\,\mathrm{dex}}$, which shows that mixing speeds are important for the production of the $p$ nuclei. This $\mathrm{O}$ shell during the merger significantly produces all $p$ nuclei except $\mathrm{{}^{74}Se}$, $\mathrm{{}^{78}Kr}$, and $\mathrm{{}^{84}Sr}$.
@@
The 3D-inspired $1\times$ scenario favours the production of the heavier $p$ nuclei compared to the MLT scenario because $\tau_{\mathrm{mix}}$ decreases as the temperature increases. Because of this, more reactions occur at the cooler temperatures where the $(n,\gamma)$ and $(\gamma,n)$ reactions are more favoured than the $(\gamma,\alpha)$ and $(\gamma,p)$ reactions. All isotopes are comparably produced to MLT or more produced except $\mathrm{{}^{74}Se}$, $\mathrm{{}^{78}Kr}$, and $\mathrm{{}^{84}Sr}$ who require the hottest temperatures for their production.
@@
As mixing speeds increase, the production increases in a non-linear and non-monotonic way. For the $50\times$ case, the average production of all $p$ nuclei is lower than the $3\times$ and $10\times$ scenarios. This is because the mixing speeds are high enough that material is advected to the bottom of the $\mathrm{O}$ shell fast enough despite the downturn, and correspondingly the lighter $p$ nuclei are generally more favoured including $\mathrm{{}^{74}Se}$, $\mathrm{{}^{78}Kr}$, and $\mathrm{{}^{84}Sr}$. Production for individual isotopes also can be non-linear and non-monotonic. For example, $\mathrm{{}^{115}Sn}$, $\mathrm{{}^{132}Ba}$, and $\mathrm{{}^{138}La}$ all increase for the $1\times$ and $3\times$ scenarios, but then their production is not as strong for the $10\times$ and $50\times$ scenarios.
@@
Another result is that isotopic pairs of the same element are not affected the same way by the downturn compared to MLT, nor by the increase in mixing speed and we find that the ratio between these isotopes is dependent on the mixing scenario. Since these isotopes are connected by $(\gamma,n)$ and $(n,\gamma)$ reactions, if the location of $D_{\alpha}=1$ for a reaction changes because of a change to the mixing speed or the presence of the decreasing diffusion profile, then the production of these isotopes will change. $\mathrm{{}^{94}Mo}$ is produced more for all mixing scenarios except for the $50\times$ scenario, where $\mathrm{{}^{92}Mo}$ has a larger $\mathrm{OP}$, and likewise $\mathrm{{}^{115}Sn}$ exhibits this behaviour when compared to $\mathrm{{}^{112}Sn}$ and $\mathrm{{}^{114}Sn}$. We find that it is possible that the ratio can tend to unity as mixing speeds increase as seen for $\mathrm{{}^{96,98}Ru}$, $\mathrm{{}^{106,108}Cd}$, and $\mathrm{{}^{112,114}Sn}$. For $\mathrm{{}^{136,138}Ce}$ the lighter isotope is always favoured as mixing speed increases and for $\mathrm{{}^{156,158}Dy}$ the opposite is true. This demonstrates the importance of the decreasing diffusion profile and increasing mixing speed is for the production of the $p$ nuclei.
III.3 Impact from varying the ingestion rate